# Which symmetry group for elementary particles with an electric charge today and in the past?

### v1

@@ Abstract
In this paper, we revisit the Kaluza-Klein theory from the perspective of the classification of elementary particles based on the coadjoint orbit method. The keystone conjecture is to consider the electric charge as an extra momentum on an equal footing with the mass and the linear momentum. We study the momentum map of the corresponding symmetry group $\hat{\mathbb{G}}_{1}$ which conserves the hyperbolic metric. We show that the electric charge is not an invariant, i.e. it depends on the reference frame, which is in contradiction with the experimental observations. In other words, it is not the symmetry group of the Universe today as we know it. To avert this paradox, we scale the fifth coordinate and consider the limit when the cylinder radius $\omega$ vanishes. For the corresponding group $\hat{\mathbb{G}}_{0}$ also of dimension 15, the charge is an invariant then independent of the frame of reference and the observer. On this ground, we propose a cosmological scenario in which the elementary particles of the early Universe are classified from the momenta of the group $\hat{\mathbb{G}}_{1}$, next the three former dimensions inflate quickly while the fifth one shrinks, leading to the 4D era in which as today the particles are characterized by the momenta of the group $\hat{\mathbb{G}}_{0}$. By this mechanism, the elementary particles can acquire electric charge as a by-product of the $4+1$ symmetry breaking of the Universe. This work opens the way to the geometric quantization of charged elementary particles. We construct the corresponding $\hat{\mathbb{G}}_{0}$-connections by pullback on the space-time. Imposing that the linear 5-momentum is parallel-transported, we recover the conservation of the charge and the equation of motion with Lorentz force. We revisit the variational relativity and obtain the field equations for both the gravitation and electromagnetic interactions with coupling terms which are negligible at the Newtonian approximation, allowing to recover Maxwell equations.
@@ 1 Introduction
Kaluza-Klein is an extension of general relativity to unify the gravitation and electromagnetism, considering a Riemannian manifold of dimension 5 by completing the 4 coordinates $(X^{1},\cdots,X^{4})$ of the space-time with an extra spacelike dimension coordinate $\hat{X}^{5}$. The pioneering paper, Kaluza [1915], started out by interpreting the 1 + 5 metric as follows [^1] Our convention in this paper is that convention that Greek indices run from 1 to 5 and Latin ones run from 1 to 4
@@
After Kaluza’s promising paper but suffering several shortcomings, another breakthrough is the work of Oskar Klein [1926a, b] who adopts also the cylinder condition but with restrictions on the coordinate changes to $X^{i}=H^{i}(X^{\prime 1},\cdots,X^{\prime 4}),\;\hat{X}^{5}=\hat{X}^{\prime 5}+h(X^{\prime 1},\cdots,X^{\prime 4})$ and then the following form the metric
@@
One of the interests of the formulation is that the electromagnetic stress-energy tensor occurs naturally at the right hand member of the first equation. Another interest is that the second equation provides Maxwell equation when the dilaton $\phi$ is set to one, but if so the last equation is satisfied only if the right hand member is null, i.e. under the too restrictive constraint that the norms of the electric and magnetic fields are equal. In contrast, if the dilaton varies, the field equations have no classical interpretation. If we continue anyway to set $\phi$ to one, the equation of geodesics reads
@@
However, the discoveries of the starting quantum mechanics suggested to Klein that the fifth dimension is curled up and microscopic. Using de Broglie wavelength $\lambda^{5}=h/p^{5}$ where $h$ is Planck constant and $p^{5}$ is the linear momentum in the fifth dimension, Klein found that the size $l_{K}$ of the space along the fifth dimension was about $10^{-30}$ cm, very small with respect to the nuclear dimension (of the order of $10^{-13}$ cm) and thereby gave an explanation for the cylinder condition in this small value. For a comparison between Kaluza and Klein approaches, Sewards [2008] can be consulted. Classical theory was completed in the 1940s and the full field equations including the scalar field were obtained almost simultaneously (Thiry [1948a, b], Jordan [1948], Scherrer [1949]).
Albert Einstein’s position with respect to Kaluza-Klein theory has changed during his scientific career. In Einstein [1920], when giving his inaugural lecture in Leyden, Einstein for the first time publicly commented positively on this unification program. However, in a letter to H. Weyl [1922], he regrets that it cannot produce non-singular rotation symmetric particle solutions. Later on, his work with W. Mayer [1931] is a a first attempt to propose an unified field theory. Commenting their work, É. Cartan [1934] interprets the torsion of the 5D space as the electromagnetic field. The last Einstein’s works on the topics are in collaboration with P. Bergmann [1938] and later on with P. Bergmann and V. Bargmann [1941]. For more details, the reader is referred to van Dongen [2000].
@@
where $C^{a}_{bc}$ are the structure constants of the group $G$. It must contain at least a $U(1)$ subgroup for the conservation of the electric charge and a $SU(2)$ subgroup for the conservation of the isospin (Kerner [1981]).
@@
Initially introduced by Bargmann [1954] to solve cohomologic difficulties to construct group actions on quantum wave functions in the Galilean setting, the Lie group that is named after him acts on a 5D space of which the fifth coordinate has the physical meaning of a specific action (in the sense of Hamilton least action principle), then it has not to be confused with Kaluza-Klein space. Bargmannian geometric structures have been studies by several authors (Duval et al. [1985, 1991], de Saxcé and Vallée [2010, 2016], Cardall [2024]). Despite of the differences with Kaluza-Klein theory, some aspects of the Bargmannian approach were a source of inspiration for this work.
@@
Another issue which has grasped researcher attention is the Kaluza-Klein cosmology. A generalized Friedmann-Lemaître-Robertson-Walker metric in 5D were introduced by Chodos and Detweiler [1980]. The idea is that while the fifth dimension has been shrinking, the other three spatial dimensions have been expanding, that could explain the large ratio of the electromagnetic to gravitational forces as a consequence of the age of the Universe, in agreement with Dirac’s large-number hypothesis. Sahdev [1984] obtained improved solutions of the field equations with perfect fluids added on the right-hand side and found that this scenario can offer a resolution to the horizon problem. However, instead of stabilizing at some small but finite value, as any reasonable physics would require, the internal radius in the fifth dimension tended to zero. Other models attempting to stabilize asymptotically the internal radius were proposed by Matzner and Mezzacappa [1985], Copeland and Toms [1985], Okada [1986], Giorgini and Kerner [1988]. These interesting works did not necessarily receive the welcome they deserved, perhaps because they were based on the theory of Kaluza-Klein whose weaknesses were known.
@@
Finally, not developed from Kaluza-Klein theory but playing a key role in this work, we have to point out Jean-Marie Souriau’s seminal work in which he proposed for the classical description of an elementary particle to study a family of homogeneous spaces for a symmetry group, the coadjoint orbits, and to classify them (Souriau [1970, 1997b]). On this ground, he introduced the geometric prequantization in Souriau [1966]. As Kaluza-Klein theory is based on an $U(1)$-principal bundle of which the base manifold is the space-time, the key idea of the prequantization is to work with a $U(1)$-principal bundle of which the base space is a symplectic manifold, a coadjoint orbit of the structure group. A tentative to combine this method with the 5D approach to treat the mass and the electric charge on an equal footing is proposed by Śniatycki [1980] but, in contrast to Kaluza-Klein theory, the fifth dimension is timelike without physical justification. The reader not familiar with this topic can consult, in addition to Souriau’s references previously cited, Kostant [1970] who is with Souriau one of the pioneers of this method, Kijowski [1977] and Woodhouse [1991].
@@
In Section 5, we apply the coadjoint orbit method to the symmetry group of Kaluza-Klein 5D space, denoted $\hat{\mathbb{G}}_{1}$, which conserves the hyperbolic metric. We show that the linear momentum along the fifth dimension, interpreted as the electric charge, is not preserved by the group, then depends on the reference frame, contradicting the observations. The aim of Section 6 is to avert this paradox by using the fact that the fifth dimension is microscopic. The result of the zoom in is to reveal a new symmetry group denoted $\hat{\mathbb{G}}_{0}$ for which the charge is invariant. Section 7 begin with the presentation of a cosmological scenario in which the particles in the early 5D Universe are classified according to the group $\hat{\mathbb{G}}_{1}$ and, after a transition phase in which the three former spatial dimensions inflate quickly while the fifth one shrinks, the particles in the 4D era are classified thanks to $\hat{\mathbb{G}}_{0}$. By this mechanism, the elementary particles can acquire electric charge as a by-product of the 4 + 1 symmetry breaking of the Universe. Next, we apply the coadjoint orbit method to the group $\hat{\mathbb{G}}_{\omega}$ for the transition between $\hat{\mathbb{G}}_{1}$ and $\hat{\mathbb{G}}_{0}$. In 5D, we generalize the spin polarisation in the form of a plane.
The second part of the trilogy is developed in Section 8. Our aim is to construct the $\hat{\mathbb{G}}_{0}$-connections on the frame principal bundle. As the zoom out leads to a singularity in the fifth dimension, we build on the space-time a pullback connection. We claim that the motion of a charged particle and the evolution of its charge are such that its linear 5-momentum is parallel-transported. The torsion free condition allows to show that the charge is conserved along the trajectory and we recover the Lorentz force.
@@ 2 Euclidean spaces
The two former subsection are simple reminder to define the terminology and notations used in the paper.
@@ 2.3 Vector product
In the particular case of the classical positive Euclidean space of dimension 3 ($n=q=3$), we use the notation $j$ instead of $\mathcal{J}$ and we recover the cross product $j(u,v)=u\times v$ and the linear map $u\mapsto j(u)$ is one-to-one from $\mathbb{R}^{3}$ into the space of $3\times 3$ skew-symmetric matrices. In the sequel, the notation $\mathcal{J}$ will be reserved to the dimension 4 and $\hat{\mathcal{J}}$ to the dimension 5. In $\mathbb{R}^{4}$ equipped with the metric of Gram’s matrix $G=diag(1,-1,-1,-1)$, the vector product of 3 vectors
@@ 2.4 Hodge operator
The adjoint [^2] To do not confuse with the adjoint of a linear map. of a $q$-form $A$ is the $(n-q)$-form *$*A$ such that
(*$(*A)(V_{1},\ldots,V_{n-q})=(-1)^{q\,(n-q)}G_{q}(A,vol(V_{1},\ldots,V_{n-q}))$
The linear map *$*:\bigwedge^{q}\mathcal{T}\rightarrow\bigwedge^{n-q}\mathcal{T}:A\mapsto*A$ is called Hodge operator. The following result can be proved:
@@
    *$**vol(V_{1},\ldots,V_{n-2})=(-1)^{n-p}\,V^{*}_{1}\wedge\ldots\wedge V^{*}_{n-2}$
@@
    *$*(V^{*}\wedge U^{*})=-vol(U,V)$
In the special case of interest : $n=4,q=2$, if $A_{M}$ is a 2-form, *$*A_{M}$ is also a 2-form identified to an skew-adjoint map denoted *$*M$. If $V^{*}_{1}\wedge V^{*}_{2}=A_{M}^{\prime}$, then $M^{\prime}=G^{-1}(V^{*}_{1}\otimes V^{*}_{2}-V^{*}_{2}\otimes V^{*}_{1})=V_{1}V^{*}_{2}-V_{2}V^{*}_{1}$. On the other hand, $\diamondsuit$ reads $A_{\mathcal{J}(V_{1},V_{2})}=*A_{M}^{\prime}=A_{*M^{\prime}}$ then:
*$*(V_{1}V^{*}_{2}-V_{2}V^{*}_{1})=\mathcal{J}(V_{1},V_{2})$ (8)
@@ 3 Coadjoint orbit method to classify the elementary particles
This Section is a simple reminder of concepts of symplectic mechanics. For more details on this topics, the reader can consult for instance Souriau [1970], Abraham and Marsden [1978], Guillemin and Sternberg [1984], Libermann and Marle [1987], Souriau [1997b]. In the sequel, all the considered Lie groups are matrix groups. Any Lie group $G$ left linearly acts on its Lie algebra $\mathfrak{g}$ by the adjoint representation
@@
Let $\left(\mathcal{M},\varpi\right)$ be a symplectic manifold, a symplectic action of a Lie group $G$
@@
The demonstration can be found for instance in Souriau [1970, 1997b, Theorem (11.34)] and [Abraham and Marsden, 1978, Section 4.2]. In the sequel, we consider only symmetry groups for which the consideration of symplectic cocycles is not relevant, then the group acts linearly.
@@
The reader can find a demonstration in Souriau [1970, 1997b, Theorem (11.34)].
@@ 4 Elementary particles in Poincaré relativity
In special relativity, the Universe $\mathcal{U}$ is represented by a 4D space [^3] In general relativity, the Universe is a manifold and we have to consider the tangent space regarding to the elementary particles.. An event $\bm{X}\in\mathcal{U}$ occuring in a reference frame at time $t$ and position $x\in\mathbb{R}^{3}$ is represented by its coordinates
@@
In the sequel, we consider only the special Lorentz transformations i.e. that belong to the connected component of the identity $\mathbb{SO}^{+}(1,3)$, a Lie group of dimension 10. The Poincaré group $\mathbb{G}$ is the set of affine transformations of $\mathbb{R}^{4}$
@@
In Souriau [1970, 1997b, Chapter 3, §14], Souriau uses the type (timelike or lightlike) of the 4-momentum vector (or energy-momentum vector) $\Pi$ for a classification of elementary systems. By these means, he obtains a large part of the physicists’ classification of elementary particles. Below, briefly summarized, his results are presented.
@@
Owing to (20), $M_{0}$ does not depends on $X$ when $X$ runs over $\mathcal{D}$. Taking into account that a non null vector orthogonal to a timelike vector is spacelike (see Souriau [1965, §28]), $I,J\in\mathbb{R}^{4}$ such that
@@
It is worth to remark that the straight line $\mathcal{D}$ is the trajectory of the particle Souriau [1970, 1997b, Chapter 3, §14]. If it is parameterized by the arc length $s$ (for the metric $G$), $U$ is the 4-velocity
@@
They are two other kind of orbits characterizing a particle without spin and a massless particle. For more details the reader is referred to Souriau [1970, 1997b, (14.24) and (14.29)].
We know that elementary particles are characterized by their mass, spin and electric charge. This last property is missing from this formalism and is usually reintroduced in the particle description as an external element to the geometric approach (see Souriau [1970, 1997b, Chapter 3, §15]). Our aim now is to develop a paradigm compatible with the experimental observations and in which the charge naturally appears as a momentum by constructing a bridge between the coadjoint orbit method and the Kaluza-Klein theory.

### v2

@@ Abstract
In this paper, we revisit the Kaluza-Klein theory from the perspective of the classification of elementary particles based on the coadjoint orbit method. The keystone conjecture is to consider the electric charge as an extra momentum on an equal footing with the mass and the linear momentum. We study the momentum map of the corresponding symmetry group $\hat{\mathbb{G}}_{1}$ which conserves the hyperbolic metric. We show that the electric charge is not an invariant, i.e. it depends on the reference frame, which is in contradiction with the experimental observations. In other words, it is not the symmetry group of the Universe today as we know it. To avert this paradox, we scale the fifth coordinate and consider the limit when the cylinder radius $\omega$ vanishes. For the corresponding group $\hat{\mathbb{G}}_{0}$, the charge is an invariant then independent of the frame of reference and the observer. On this ground, we propose a cosmological scenario in which the elementary particles of the early Universe are classified from the momenta of the group $\hat{\mathbb{G}}_{1}$, next the three former dimensions inflate quickly while the fifth one shrinks, leading to the 4D era in which as today the particles are characterized by the momenta of the group $\hat{\mathbb{G}}_{0}$. By this mechanism, the elementary particles can acquire electric charge as a by-product of the $4+1$ symmetry breaking of the Universe. This work opens the way to the geometric quantization of charged elementary particles. We construct the corresponding $\hat{\mathbb{G}}_{0}$-connections by pullback on the space-time. Imposing that the linear 5-momentum is parallel-transported, we recover the conservation of the charge and the equation of motion with Lorentz force. We revisit the variational relativity and obtain the field equations for both the gravitation and electromagnetic interactions with coupling terms which are negligible at the Newtonian approximation, allowing to recover Maxwell equations.
@@ 1 Introduction
Kaluza-Klein theory is an extension of general relativity to unify the gravitation and electromagnetism, considering a Riemannian manifold of dimension 5 by completing the 4 coordinates $(X^{1},\cdots,X^{4})$ of the space-time with an extra spacelike dimension coordinate $\hat{X}^{5}$.
Theodor Kaluza’s pioneering paper, [23], started out by interpreting the 1 + 5 metric as follows [^1] Our convention in this paper is that convention that Greek indices run from 1 to 5 and Latin ones run from 1 to 4
@@
After Kaluza’s promising paper but suffering several shortcomings, another breakthrough is the work of Oskar Klein [28, 29] who adopts also the cylinder condition but with restrictions on the coordinate changes to $X^{i}=H^{i}(X^{\prime 1},\cdots,X^{\prime 4}),\;\hat{X}^{5}=\hat{X}^{\prime 5}+h(X^{\prime 1},\cdots,X^{\prime 4})$ and then the following form the metric
@@
One of the interests of the formulation is that the electromagnetic stress-energy tensor occurs naturally at the right hand member of the first equation. Another interest is that the second equation provides Maxwell equation when the dilaton $\phi$ is set to one, but if so the last equation is satisfied only if the right hand member is null, i.e. under the too restrictive constraint that the norms of the electric and magnetic fields are equal. In contrast, if the dilaton varies, the field equation has no classical interpretation. If we continue anyway to set $\phi$ to one, the equation of geodesics reads
@@
However, the discoveries of the starting quantum mechanics suggested to Klein that the fifth dimension is curled up and microscopic. Using de Broglie wavelength $\lambda^{5}=h/p^{5}$ where $h$ is Planck constant and $p^{5}$ is the linear momentum in the fifth dimension, Klein found that the size $l_{K}$ of the space along the fifth dimension was about $10^{-31}$ cm, very small with respect to the nuclear dimension (of the order of $10^{-13}$ cm) and thereby gave an explanation for the cylinder condition in this small value. For a comparison between Kaluza and Klein approaches, [42] can be consulted. Classical theory was completed in the 1940s and the full field equations including the scalar field were obtained almost simultaneously ([51, 52, 22, 41]).
Albert Einstein’s position with respect to Kaluza-Klein theory has changed during his scientific career. In [15], when giving his inaugural lecture in Leyden, Einstein for the first time publicly commented positively on this unification program. However, in a [16], he regrets that it cannot produce non-singular rotation symmetric particle solutions. Later on, his work with W. Mayer [17] is a a first attempt to propose an unified field theory. Commenting their work, É. Cartan [6] interprets the torsion of the 5D space as the electromagnetic field. The last Einstein’s works on the topics are in collaboration with P. Bergmann [18] and later on with P. Bergmann and V. Bargmann [19]. For more details, the reader is referred to [53].
@@
where $C^{a}_{bc}$ are the structure constants of the group $G$. It must contain at least a $U(1)$ subgroup for the conservation of the electric charge and a $SU(2)$ subgroup for the conservation of the isospin [25].
@@
Initially introduced by V. Bargmann in [2] to solve cohomologic difficulties to construct group actions on quantum wave functions in the Galilean setting, the Lie group that is named after him acts on a 5D space of which the fifth coordinate has the physical meaning of a specific action (in the sense of Hamilton least action principle), then it has not to be confused with Kaluza-Klein space. Bargmannian geometric structures have been studies by several authors ([12, 13, 9, 10, 5]). Despite of the differences with Kaluza-Klein theory, some aspects of the Bargmannian approach were a source of inspiration for this work.
@@
Another issue which has grasped researcher attention is the Kaluza-Klein cosmology. A generalized Friedmann-Lemaître-Robertson-Walker metric in 5D were introduced in [7]. The idea is that while the fifth dimension has been shrinking, the other three spatial dimensions have been expanding, that could explain the large ratio of the electromagnetic to gravitational forces as a consequence of the age of the Universe, in agreement with Dirac’s large-number hypothesis. In [40], Sahdev obtained improved solutions of the field equations with perfect fluids added on the right-hand side and found that this scenario can offer a resolution to the horizon problem. However, instead of stabilizing at some small but finite value, as any reasonable physics would require, the internal radius in the fifth dimension tended to zero. Other models attempting to stabilize asymptotically the internal radius were proposed in [36, 8, 37, 26]. These interesting works did not necessarily receive the welcome they deserved, perhaps because they were based on the theory of Kaluza-Klein whose weaknesses were known.
@@
Finally, not developed from Kaluza-Klein theory but playing a key role in this work, we have to point out Jean-Marie Souriau’s seminal work in which he proposed for the classical description of an elementary particle to study a family of homogeneous spaces for a symmetry group, the coadjoint orbits, and to classify them ([48, 50]). On this ground, he introduced the geometric prequantization in [47]. As Kaluza-Klein theory is based on an $U(1)$-principal bundle of which the base manifold is the space-time, the key idea of the prequantization is to work with a $U(1)$-principal bundle of which the base space is a symplectic manifold, a coadjoint orbit of the structure group. A tentative to combine this method with the 5D approach to treat the mass and the electric charge on an equal footing is proposed in [43] but, in contrast to Kaluza-Klein theory, the fifth dimension is timelike without physical justification. The reader not familiar with this topic can consult, in addition to Souriau’s references previously cited, Kostant [32] who is with Souriau one of the pioneers of this method, [27] and [54].
@@
In Section 5, we apply the coadjoint orbit method to the symmetry group of Kaluza-Klein 5D space, denoted $\hat{\mathbb{G}}_{1}$, which conserves the hyperbolic metric. We show that the linear momentum along the fifth dimension, interpreted as the electric charge, is not preserved by the group, then depends on the reference frame, contradicting the observations. The aim of Section 6 is to avert this paradox by using the fact that the fifth dimension is overwhelmingly small. The result of the zoom in is to reveal a new symmetry group denoted $\hat{\mathbb{G}}_{0}$ for which the charge is invariant. Section 7 begins with the presentation of a cosmological scenario in which the particles in the early 5D Universe are classified according to the group $\hat{\mathbb{G}}_{1}$ and, after a transition phase in which the three former spatial dimensions inflate quickly while the fifth one shrinks, the particles in the 4D era are classified thanks to $\hat{\mathbb{G}}_{0}$. By this mechanism, the elementary particles can acquire electric charge as a by-product of the 4 + 1 symmetry breaking of the Universe. Next, we apply the coadjoint orbit method to the group $\hat{\mathbb{G}}_{\omega}$ for the transition between $\hat{\mathbb{G}}_{1}$ and $\hat{\mathbb{G}}_{0}$. In 5D, we generalize the spin polarisation in the form of a plane.
The second part of the trilogy is developed in Section 8. Our aim is to construct $\hat{\mathbb{G}}_{0}$-connections on the frame principal bundle. As the zoom out leads to a singularity in the fifth dimension, we build on the space-time a pullback connection. We claim that the motion of a charged particle and the evolution of its charge are such that its linear 5-momentum is parallel-transported. The torsion free condition allows to show that the charge is conserved along the trajectory and we recover the Lorentz force.
@@ 2 Euclidean spaces
The two former subsection are simple reminders to define the terminology and notations used in the paper.
@@ 2.3 Vector product
In the particular case of the classical positive Euclidean space of dimension 3 ($n=p=3$), we use the notation $j$ instead of $\mathcal{J}$ and we recover the cross product $j(u,v)=u\times v$ and the linear map $u\mapsto j(u)$ is one-to-one from $\mathbb{R}^{3}$ into the space of $3\times 3$ skew-symmetric matrices. In the sequel, the notation $\mathcal{J}$ will be reserved to the dimension 4 and $\hat{\mathcal{J}}$ to the dimension 5. In $\mathbb{R}^{4}$ equipped with the metric of Gram’s matrix $G=diag(1,-1,-1,-1)$, the vector product of 3 vectors
@@ 2.4 Hodge operator
The adjoint [^2] To do not confuse with the adjoint of a linear map. of a $q$-form $A$ is the $(n-q)$-form $*A$ such that
$(*A)(V_{1},\ldots,V_{n-q})=(-1)^{q\,(n-q)}G_{q}(A,vol(V_{1},\ldots,V_{n-q}))$
The linear map $*:\bigwedge^{q}\mathcal{T}\rightarrow\bigwedge^{n-q}\mathcal{T}:A\mapsto*A$ is called Hodge operator. The following result can be proved:
@@
    $*vol(V_{1},\ldots,V_{n-2})=(-1)^{n-p}\,V^{*}_{1}\wedge\ldots\wedge V^{*}_{n-2}$
@@
    $*(V^{*}\wedge U^{*})=-vol(U,V)$
In the special case of interest : $n=4,q=2$, if $A_{M}$ is a 2-form, $*A_{M}$ is also a 2-form identified to an skew-adjoint map denoted $*M$. If $V^{*}_{1}\wedge V^{*}_{2}=A_{M}^{\prime}$, then $M^{\prime}=G^{-1}(V^{*}_{1}\otimes V^{*}_{2}-V^{*}_{2}\otimes V^{*}_{1})=V_{1}V^{*}_{2}-V_{2}V^{*}_{1}$. On the other hand, $\diamondsuit$ reads $A_{\mathcal{J}(V_{1},V_{2})}=*A_{M}^{\prime}=A_{*M^{\prime}}$ then:
$*(V_{1}V^{*}_{2}-V_{2}V^{*}_{1})=\mathcal{J}(V_{1},V_{2})$ (8)
@@ 3 Coadjoint orbit method to classify the elementary particles
This Section is a simple reminder of concepts of symplectic mechanics. For more details on this topics, the reader can consult for instance [48, 1, 20, 35, 50]. In the sequel, all the considered Lie groups are matrix groups. Any Lie group $G$ left linearly acts on its Lie algebra $\mathfrak{g}$ by the adjoint representation
@@
Let $\left(\mathcal{M},\varpi\right)$ be a symplectic manifold, an action of a Lie group $G$ on $\mathcal{M}$
@@
The demonstration can be found for instance in Theorem (11.17) of [48, 50] and Section 4.2 of [1]. In the sequel, we consider only symmetry groups for which the consideration of symplectic cocycles is not relevant, then the group acts linearly.
@@
The reader can find a demonstration in Theorem (11.34) of [48, 50].
@@ 4 Elementary particles in Poincaré relativity
In special relativity, the Universe $\mathcal{U}$ is represented by a 4D hyperbolic Euclidean space [^3] In general relativity, the Universe is a manifold and we have to consider the tangent space regarding to the elementary particles.. An event $\bm{X}\in\mathcal{U}$ occuring in a reference frame at time $t$ and position $x\in\mathbb{R}^{3}$ is represented by its coordinates
@@
In the sequel, we consider only the special Lorentz transformations i.e. that belong to the connected component of the identity $\mathbb{SO}^{+}(1,3)$, a Lie group of dimension 6. The Poincaré group $\mathbb{G}$ is the set of affine transformations of $\mathbb{R}^{4}$
@@
In Chapter 3, §14 of [48, 50], Souriau uses the type (timelike or lightlike) of the 4-momentum vector (or energy-momentum vector) $\Pi$ for a classification of elementary systems. By these means, he obtains a large part of the physicists’ classification of elementary particles. Below, briefly summarized, his results are presented.
@@
Owing to (20), $M_{0}$ does not depends on $X$ when $X$ runs over $\mathcal{D}$. Taking into account that a non null vector orthogonal to a timelike vector is spacelike (see §28 of [46]), we introduce $I,J\in\mathbb{R}^{4}$ such that
@@
It is worth to remark that the straight line $\mathcal{D}$ is the trajectory of the particle (Chapter 3, §14 of [48, 50]). If it is parameterized by the arc length $s$ (for the metric $G$), $U$ is the 4-velocity
@@
They are two other kind of orbits characterizing a particle without spin and a massless particle. For more details the reader is referred to (14.24) and (14.29) of [48, 50].
We know that elementary particles are characterized by their mass, spin and electric charge. This last property is missing from this formalism and is usually reintroduced in the particle description as an external element to the geometric approach (see Chapter 3, §15 of [48, 50]). Our aim now is to develop a paradigm compatible with the experimental observations and in which the charge naturally appears as a momentum by constructing a bridge between the coadjoint orbit method and the Kaluza-Klein theory.